OpenStudy (anonymous):

HELP! 1) A right triangle has one leg of length 8 and another leg of length 9. What is the length of the hypotenuse of the triangle? 2)A right triangle has a hypotenuse of length 15 and a leg of 9. what is the length of the other leg of the triangle?

OpenStudy (andriod09):

try this equation: a^2+b^2=c^2

OpenStudy (andriod09):

8^2+9^2=c^2 15^2+9^2=c62

OpenStudy (anonymous):

OK AND THEN?

OpenStudy (anonymous):

225+81= c^2?

OpenStudy (andriod09):

the you square it.

OpenStudy (anonymous):

how???

OpenStudy (andriod09):

using square root, you basicly get rid of the squared "c". i,e it would be 15+9=c

OpenStudy (anonymous):

wait but the hypotenuse is 15 isnt that suppose to be C?

OpenStudy (andriod09):

the the hypotenuse is 15+9 not 15

OpenStudy (andriod09):

thats problem #2 also

OpenStudy (anonymous):

wait i am confused lol ok you use the equation a^2+b^2=c^2 right? like the hypotenuse is a^2? not c^2?

OpenStudy (andriod09):

no, the hypotenuse is a^2_b^2, but you took away the squares remember? so then it would be 15+9=c

OpenStudy (anonymous):

okkkk and then? i just add it?

OpenStudy (andriod09):

yea, the answe rfor #2 would be 15+9=c

hartnn (hartnn):

nopes, it won't be c=15+9.

OpenStudy (andriod09):

yes it would hartnn

OpenStudy (anonymous):

omg.........what??? i am so confused now lol

hartnn (hartnn):

for 1st problem: hypotenuse = c \(a^2+b^2=c^2 \\ 8^2+9^2=c^2=145 \\ c=\sqrt{145}=12.04\)

OpenStudy (anonymous):

i got that one!!! :D but how do i solve the other one?

OpenStudy (anonymous):

dont i have to do the same thing?

OpenStudy (andriod09):

okay, hartnn, you just confused me, and im in algebre 3. what the hell did you just do??

hartnn (hartnn):

for 2nd problem: hypotenuse c = 15 . \(a^2+b^2=c^2 \\ 9^2+b^2=15^2 \\ b^2=15^2-9^2 \\ b^2=225-81=144 \\b=\sqrt{144} \\b=12\)

hartnn (hartnn):

@andriod09 which step u are confused with ?

hartnn (hartnn):

@ayojenn got the 2nd one ?

OpenStudy (andriod09):

everything! 1st, i cant see equations used with the equation button, so thats one, and 2, thats not how i was taught to to pythagorium theory.

OpenStudy (anonymous):

yesss that what i was trying to tell @andriod09 , i knew the hypotenuse was c^2

OpenStudy (andriod09):

ik that, thats what i was getting to, ill be right back.

hartnn (hartnn):

the hypotenuse = c and not c^2

OpenStudy (anonymous):

wait what???? where did you get 82?

hartnn (hartnn):

for 2nd problem: hypotenuse c = 15 . a2+b2=c2 9^2+b^2=15^2 b2=15^2−9^2 b2=225−81=144b=144 b=12 i never got 82

OpenStudy (anonymous):

hartnn Best Response 0 ohh....wait 1st problem: a2+b2=c2 82+92=c2=145 <--- 82 c=√145=12.04

hartnn (hartnn):

ohh....wait 1st problem: a^2+b^2=c^2 8^2+9^2=c^2=145 c=√145=12.04

hartnn (hartnn):

i meant that^ sorry for confusion.....

OpenStudy (anonymous):

yes thats what i got :D lol its okk

OpenStudy (anonymous):

ok theres another problem but i solved it just let me know if its correct ok?

hartnn (hartnn):

okay :)

OpenStudy (anonymous):

A right triangles has a hypotenuse of length 21 and a leg of length 14. What is the length of the other leg of the triangle

OpenStudy (anonymous):

this is what i got...

OpenStudy (anonymous):

14^2+v^2=21^2 b=21^2-14^2 441-196 square root 245= 15.65?

hartnn (hartnn):

absolutely correct! :) u are good at this :D

OpenStudy (anonymous):

omg yay i learned!!! haha thank you :D

hartnn (hartnn):

welcome :)